# From Chat Logs to Collective Insights: Aggregative Question Answering

### v1

@@ 4.5 Ablation Studies
We conducted ablation studies on a stratified 10% subset of the benchmark, selected based on condition and target types.
Retrieval Effectiveness is Critical
Retrieval performance has a significant impact on the final results. Table 5 presents the performance of o4-mini under varying recall rates achieved by different retrieval methods. We observe that higher recall rates consistently lead to improved NDCG scores.
@@
We compared various retrieval approaches, including vector-based embeddings, BM25, random, and ground-truth retrieval. Figure 5 shows recall rates for different retrieval strategies. PROBE consistently provided substantial improvements over standard RAG, with the highest recall from PROBE-Dense (summarized). Removing either the generated query or filtering steps notably degraded PROBE’s retrieval effectiveness (Table 4).
Current models lack effective aggregational reasoning
To explore model capabilities under ideal conditions, we performed experiments using oracle (ground-truth) documents as context (Table 6). All models performed better when given summarized contexts rather than raw conversations, indicating challenges in aggregating information from longer, noisier texts.
@@
We further analyzed how performance varied with the number of provided conversations (Figure 6). Weaker models such as Gemma3 and Qwen3 showed a substantial performance gap between raw and summarized contexts, even when given the same number of conversations, highlighting their limited ability to implicitly extract relevant information. Stronger models like GPT-4.1-mini and o4-mini showed a smaller initial gap, but this gap widened notably when the context increased to 100 documents, demonstrating that even advanced models struggle with aggregating and reasoning effectively over extensive raw contexts.
@@
Unlike standard RAG tasks, Aggregative Question Answering fundamentally relies on a broader set of documents. As more documents were provided, models improved significantly in answering aggregative questions (Figure 7). This finding validates that aggregative question answering requires extensive context and global dataset knowledge.
@@
We evaluated Qwen3-32B (with optional reasoning via “think” mode) to measure the effect of explicit reasoning. Results (Table 7) consistently showed reasoning led to significant performance improvements across all experimental setups, indicating aggregative question answering demands substantial reasoning abilities.
@@ 5 Future Research Directions
In this work, we experimented with several reasoning-capable models and observed that current models typically have limited context windows, and performance degrades sharply as the length of the input context increases. Developing efficient and accurate methods for reasoning over very long textual contexts remains an important open problem.
@@
In real-world scenarios, chatbot conversations often arrive in continuous streams rather than static collections. Future research could explore methods to dynamically update aggregational insights as new interactions occur in real time. Ideally, conversational agents would continuously integrate information from ongoing interactions—similar to how humans update their understanding based on new experiences—to maintain up-to-date and adaptive aggregational knowledge.
@@ 6 Conclusion
In this paper, we introduced Aggregative Question Answering, a new task aimed at extracting collective insights from large-scale conversational data generated by interactions between users and LLM-powered chatbots. To facilitate research in this area, we constructed the WildChat-AQA benchmark, comprising 6,027 aggregational questions derived from 182,330 real-world chatbot conversations. Our experiments demonstrate that existing state-of-the-art methods, including fine-tuning, retrieval-augmented generation (RAG), and even an improved RAG approach specifically adapted for this task—PROBE (Probing Retrieval Of Broad Evidence)—struggle significantly, either failing to reason effectively at the necessary global scale or incurring prohibitively high computational costs. Looking ahead, we believe addressing these challenges would enable future models to better derive meaningful user and societal insights from large-scale conversational data.
@@
Although we employed powerful large language models and pipelines (such as GPT-4o and TnT-LLM) to infer attributes and assign taxonomy labels, errors and inconsistencies may still occur due to model hallucinations or instruction misalignment. Additionally, real-world conversational data is inherently noisy, ambiguous, and challenging to categorize neatly, making it impossible to ensure completely error-free annotations.
@@
Aggregative questions in WildChat-AQA were generated by prompting GPT-4.1 to translate structured database queries into natural-language questions. While effective, this method may result in questions that still feel somewhat templated or artificial, potentially limiting their naturalness and realism compared to genuinely human-authored queries.
@@
Aggregative Question Answering opens promising avenues for impactful real-world analytics but also raises potential ethical and societal concerns, particularly when insights relate to sensitive topics such as elections, public opinion, or public health—areas that could potentially be susceptible to manipulation. To mitigate such risks, all experiments conducted in this work rely exclusively on the publicly available and anonymized WildChat dataset, which is explicitly intended for open research purposes (licensed under ODC-BY). By introducing WildChat-AQA as an open benchmark, we aim to empower transparent academic research that responsibly explores both the capabilities and risks associated with aggregational analytics. Our goal is to encourage the open research community to better understand and evaluate these powerful systems, rather than relying solely on proprietary analyses conducted behind closed doors.
@@ B.2 Language Distribution
We provide a statistics of all language involved in the conversations in Table 9.
@@ B.3 Keywords Cloud
To illustrate the result of keywords categorization, we build a keywords cloud in Figure 8
@@ Appendix C Data Visualization Demonstration
We developed an interactive data visualization interface using React.js and Next.js for the frontend, and FastAPI for the backend implementation. MongoDB serves as the database system. An overview of the interface is shown in Figure 9. Users can filter generated questions using a configurable question filter, as illustrated in Figure 10.
@@ Appendix D Experiment Implementation Details
For inference with open-source models, we utilized vLLM v0.8.5.post1. The sampling hyperparameters used during inference are detailed in Table 11.
@@
For query generation in PROBE, we use GPT-4.1-mini as query and filter generator with top_p = 0.5 and top_k = 0.5.
@@ E.2 LLM-based keywords and summarization extraction
To perform TnT-LLM for topic discovery, we begin by extracting keywords and summaries from raw conversations. Specifically, we prompt GPT-4o to generate both the keyword set and a concise summarization of each conversation. The extracted keywords span a diverse set of semantic types, including persons, technologies, scientific terms, foods, demographic terms, organizations, locations, events, artworks, programming languages, product brands, and financial terms. The complete prompt used for this extraction process is shown in Figure 14.
@@ E.3.1 Topic Discovery and Assignment
Subsequently, we apply the topic discovery algorithm detailed in Algorithm 1. The initial taxonomy generated is visualized in Figure 15, while the prompt used for topic refinement is shown in Figure 16. For all topic discovery steps, we employ GPT-4o as the underlying language model, using hyperparameters and . To perform efficient KMeans clustering, we utilize the FAISS library Douze et al. (2025). Unlike the original TnT-LLM method, which relies on LLMs for taxonomy refinement, we manually resolve conflicts and enforce mutual exclusivity among the discovered topics.
@@
Using the generated topics and corresponding taxonomy, we assign a topic ID to each conversation. This assignment process can be formulated as a multi-label classification task. The labeling is performed by GPT-4o using the assignment prompt illustrated in Figure 17. The prompt is carefully designed to mitigate common errors identified through a manual inspection of a small validation set consisting of 400 examples.
@@ E.3.2 Subtopic Discovery and Assignment
Prompt GPT-4o to check the result of topic assignment and summarize the raw conversation from the perspective of major topic using the prompt shown in Figure 18.
@@
For each target number of subtopics ,we execute Algorithm 1 with parameters using topic-specific initial and update prompts as illustrated in Figure 19 and Figure 20. The model used for subtopic discovery is OpenAI-o1, selected for its strong reasoning capabilities. To enforce the desired number of generated subtopics at the start of the iteration, we replace the placeholder “{min_class_number_requirement}” in Figure 19 with instruction “- Generate NO LESS THAN topics.” .
@@
After generating the taxonomy for each , we randomly sample 10% of data instances from the current topic—capped at a maximum of 1000 samples. We then query the o3-mini model, which has strong reasoning ability, using the prompt provided in Figure 21. This yields a set of predicted labels , along with corresponding relevance scores between 0-10, each ranging from 0 to 10. We then compute a quality score for each generated taxonomy using the following equations:
@@
Finally, we label all data samples using the prompt illustrated in Figure 21, with the o3-mini model. For each topic, we select the best-performing taxonomy and use it to annotate all corresponding samples.
@@ E.4 Topic Label Quality Control
After completing the labeling pipeline, we still observed some false positives upon manual inspection. To address this, we conducted an additional verification step—similar to the initial phase of the subtopic discovery pipeline—by reviewing each data sample alongside its raw conversation, assigned label, and label description, using the o3-mini model and the prompt shown in Figure 22. Following this verification, we removed all samples that lacked a valid label assignment or were assigned the Undefined label at either the topic or subtopic level. This filtering ensured that the final dataset aligned with the discovered taxonomy, ultimately reducing the dataset size to approximately 182k examples.
@@ E.5.1 LLM Based Aggregation
Assuming that the same word used by the same user conveys a consistent meaning, we first associate each user’s keyword with its corresponding description, extracted at the beginning of the process. We then employ o3-mini to cluster these raw keywords into semantically coherent groups, corresponding to categories including “Programming Language”, “Video Games”, “Tabletop Games”, “Manga/Anime”, “Film”, “TV Show”, “Western Cartoon/Comic”, “Book”, “Musical”, and “Public Figure” , using the prompt illustrated in Figure 23.
@@ E.6 Question Proposal
We generate questions through a brute-force search over various combinations and quantities of conditions. The full set of considered conditions is shown in Table 1. Specifically, we enumerate all possible attribute combinations containing 0 to 3 conditions and manually select 73 meaningful combinations that can be naturally expressed in language. The selected combinations are listed in Table 8.
@@ E.7 Question Generation
Given a set of condition types, corresponding values, and a target value, we prompt GPT-4.1 to generate natural language questions using the template shown in Figure 24.
@@ E.8 Question Quality Control
We employ o4-mini for final quality control. Specifically, o4-mini is used to rank target candidates under two settings: (1) without any supporting context, and (2) with supporting context provided in the form of either summaries or raw conversations, using the prompting format shown in Figure 25.For each instance, we compute the instance-wise NDCG@10 score in the no-context setting, denoted as , and define the contextual score as , where and are scores under raw and summarized contexts, respectively.

### v2

@@ 4.5 Ablation Studies
We conduct ablation studies on a stratified 10% subset of the benchmark, selected based on the condition and target types.
Retrieval effectiveness is crucial.
Retrieval quality substantially affects final performance. Table˜5 reports the results of o4-mini under varying recall rates from different retrieval methods. Higher recall rates consistently yield better NDCG scores.
@@
We compare various retrieval approaches, including vector-based embeddings, BM25, random, and ground-truth retrieval. Figure˜5 shows recall rates for different retrieval strategies. PROBE consistently provided substantial improvements over standard RAG, with the highest recall from PROBE-Dense (summarized). Removing either the generated query or filtering steps notably degrades PROBE’s retrieval effectiveness as shown in Table˜4.
Existing models lack effective aggregational reasoning capabilities.
To evaluate model capabilities under ideal conditions, we perform experiments using oracle documents as context. Table˜6 shows that all models perform better when given summarized contexts rather than raw conversations, indicating challenges in aggregating information from longer, noisier texts.
@@
We further analyze how performance varies with the number of provided conversations (Figure˜6). Weaker models such as Gemma3 and Qwen3 show a substantial performance gap between raw and summarized contexts, even when given the same number of conversations, highlighting their limited ability to implicitly extract relevant information. Stronger models like GPT-4.1-mini and o4-mini show a smaller initial gap, but this gap widens notably when the context is extended to 100 documents, demonstrating that even advanced models struggle with aggregating and reasoning effectively over extensive raw contexts.
@@
Unlike standard RAG tasks, Aggregative Question Answering fundamentally relies on a broader set of documents. As more documents are provided, models improve significantly in answering aggregative questions (Figure˜7). This finding validates that aggregative question answering requires extensive context and global dataset knowledge.
@@
We evaluate Qwen3-32B with the “think” mode on to measure the effect of explicit reasoning. The results (see Table˜7) consistently show reasoning led to significant performance improvements across all experimental setups, indicating aggregative question answering demands substantial reasoning abilities.
@@ 5 Future Research Directions
In this work, we experiment with several reasoning-capable models and observe that current models typically have limited context windows, and performance degrades sharply as the length of the input context increases. Developing efficient and accurate methods for reasoning over very long textual contexts remains an important open problem.
@@
In real-world scenarios, chatbot conversations often arrive in continuous streams rather than static collections. Future research could explore methods to dynamically update aggregational insights as new interactions occur in real time. Ideally, conversational agents would continuously integrate information from ongoing interactions, similar to how humans update their understanding based on new experiences, to maintain up-to-date and adaptive aggregational knowledge.
@@ 6 Conclusion
In this paper, we introduce Aggregative Question Answering, a new task aimed at extracting collective insights from large-scale conversational data generated by interactions between users and LLM-powered chatbots. To facilitate research in this area, we construct the WildChat-AQA benchmark, comprising 6,027 aggregational questions derived from 182,330 real-world chatbot conversations. Our experiments demonstrate that existing state-of-the-art methods, including fine-tuning, Retrieval-Augmented Generation (RAG), and even an improved RAG approach specifically adapted for this task struggle significantly, either failing to reason effectively at the necessary global scale or incurring prohibitively high computational costs. Looking ahead, we believe that addressing these challenges would enable future models to better derive meaningful user and societal insights from large-scale conversational data.
@@
Although we employ powerful large LLMs and pipelines such as GPT-4o and TnT-LLM to infer attributes and assign taxonomy labels, errors and inconsistencies may occur due to model hallucinations or instruction misalignment. Specifically, hallucinations might affect both the inferred topics (summaries used to construct taxonomies) and the extracted keywords, potentially introducing noise or inaccuracies into the benchmark. To quantify these potential errors, we conduct human evaluations measuring the agreement between human annotations and LLM annotations for both topic extraction (Table 2) and keyword extraction. Although these evaluations confirm moderate to high accuracy, we acknowledge that some errors remain inevitable. Additionally, real-world conversational data are inherently noisy, ambiguous, and challenging to categorize neatly, making completely error-free annotations unattainable. We encourage future users of our dataset to remain aware of these limitations when interpreting experimental results.
@@
Aggregative questions in WildChat-AQA were generated by prompting GPT-4.1 to translate structured database queries into natural-language questions. While effective and typically resulting in simple and straightforward queries, this method may introduce stylistic, syntactic, and semantic artifacts. Models trained on our data can potentially overfit to the stylistic patterns of LLM-generated questions, which could limit the validity of the introduced benchmark. Consequently, strong performance on WildChat-AQA may not directly generalize to success on genuinely human-authored aggregative questions, which tend to be linguistically richer and more diverse. We thus consider strong performance on our benchmark as a necessary but not sufficient condition for aggregative question-answering capabilities in real-world scenarios.
@@
Aggregative Question Answering opens promising avenues for real-world analytics but also raises potential ethical and societal concerns, particularly when insights relate to sensitive topics such as elections, public opinion, or public health—areas that could potentially be susceptible to manipulation. To reduce the risks of reinforcing stereotypes or enabling sensitive demographic profiling, we avoided constructing questions targeting protected attributes. Moreover, all experiments conducted in this work rely exclusively on the publicly available and anonymized WildChat dataset, which is explicitly intended for open research purposes (licensed under ODC-BY). By introducing WildChat-AQA as an open benchmark, we aim to empower transparent academic research that responsibly explores both the capabilities and risks associated with aggregational analytics. Our goal is to encourage the open research community to evaluate these powerful systems, rather than relying solely on proprietary analyses conducted behind closed doors.
Acknowledgements
Yuntian Deng acknowledges support from an NSERC Discovery Grant (RGPIN-2024-05178), a Starter Grant from the University of Waterloo, and research funding provided by Manulife through the Waterloo Data and Artificial Intelligence Institute. We also thank OpenAI’s Research Access Program for providing API credits. Wentao Zhang is supported in part by these sources and by the Dr. Derick Wood Graduate Scholarship, generously funded by Ms. Mary Chen.
@@ B.2 Language Distribution
We provide a statistics of all language involved in the conversations in Table˜9.
@@ B.3 Keywords Cloud
To illustrate the result of keywords categorization, we build a keywords cloud in Figure˜8.
@@ Appendix C Data Visualization Demonstration
We developed an interactive data visualization interface using React.js and Next.js for the frontend, and FastAPI for the backend implementation. MongoDB serves as the database system. An overview of the interface is shown in Figure˜9. Users can filter generated questions using a configurable question filter, as illustrated in Figure˜10.
@@ Appendix D Experiment Implementation Details
For text pre-processing of RAG, we chunked raw conversations into segments of at most 512 tokens with a maximum overlap of 128 tokens, preserving sentence boundaries wherever possible. For summarized conversations, no chunking is performed due to relatively short text length.
For inference with open-source models, we utilized vLLM v0.8.5.post1. The sampling hyperparameters used during inference are detailed in Table˜11.
@@
For broad query generation in PROBE, we use GPT-4.1-mini as query and filter generator with top_p = 0.5 and top_k = 0.5.
@@ E.2 LLM-based keywords and summarization extraction
To perform TnT-LLM for topic discovery, we begin by extracting keywords and summaries from raw conversations. Specifically, we prompt GPT-4o to generate both the keyword set and a concise summarization of each conversation. The extracted keywords span a diverse set of semantic types, including persons, technologies, scientific terms, foods, demographic terms, organizations, locations, events, artworks, programming languages, product brands, and financial terms. The complete prompt used for this extraction process is shown in Figure˜14.
@@ E.3.1 Topic Discovery and Assignment
Subsequently, we apply the topic discovery algorithm detailed in Algorithm˜1. The initial taxonomy generated is visualized in Figure˜15, while the prompt used for topic refinement is shown in Figure˜16. For all topic discovery steps, we employ GPT-4o as the underlying language model, using hyperparameters and . To perform efficient KMeans clustering, we utilize the FAISS library Douze et al. (2025). Unlike the original TnT-LLM method, which relies on LLMs for taxonomy refinement, we manually resolve conflicts and enforce mutual exclusivity among the discovered topics.
@@
Using the generated topics and corresponding taxonomy, we assign a topic ID to each conversation. This assignment process can be formulated as a multi-label classification task. The labeling is performed by GPT-4o using the assignment prompt illustrated in Figure˜17. The prompt is carefully designed to mitigate common errors identified through a manual inspection of a small validation set consisting of 400 examples.
@@ E.3.2 Subtopic Discovery and Assignment
Prompt GPT-4o to check the result of topic assignment and summarize the raw conversation from the perspective of major topic using the prompt shown in Figure˜18.
@@
For each target number of subtopics ,we execute Algorithm˜1 with parameters using topic-specific initial and update prompts as illustrated in Figure˜19 and Figure˜20. The model used for subtopic discovery is OpenAI-o1, selected for its strong reasoning capabilities. To enforce the desired number of generated subtopics at the start of the iteration, we replace the placeholder “{min_class_number_requirement}” in Figure˜19 with instruction “- Generate NO LESS THAN topics.” .
@@
After generating the taxonomy for each , we randomly sample 10% of data instances from the current topic—capped at a maximum of 1000 samples. We then query the o3-mini model, which has strong reasoning ability, using the prompt provided in Figure˜21. This yields a set of predicted labels , along with corresponding relevance scores between 0-10, each ranging from 0 to 10. We then compute a quality score for each generated taxonomy using the following equations:
@@
Finally, we label all data samples using the prompt illustrated in Figure˜21, with the o3-mini model. For each topic, we select the best-performing taxonomy and use it to annotate all corresponding samples.
@@ E.4 Topic Label Quality Control
After completing the labeling pipeline, we still observed some false positives upon manual inspection. To address this, we conducted an additional verification step—similar to the initial phase of the subtopic discovery pipeline—by reviewing each data sample alongside its raw conversation, assigned label, and label description, using the o3-mini model and the prompt shown in Figure˜22. Following this verification, we removed all samples that lacked a valid label assignment or were assigned the Undefined label at either the topic or subtopic level. This filtering ensured that the final dataset aligned with the discovered taxonomy, ultimately reducing the dataset size to approximately 182k examples.
@@ E.5.1 LLM Based Aggregation
Assuming that the same word used by the same user conveys a consistent meaning, we first associate each user’s keyword with its corresponding description, extracted at the beginning of the process. We then employ o3-mini to cluster these raw keywords into semantically coherent groups, corresponding to categories including “Programming Language”, “Video Games”, “Tabletop Games”, “Manga/Anime”, “Film”, “TV Show”, “Western Cartoon/Comic”, “Book”, “Musical”, and “Public Figure” , using the prompt illustrated in Figure˜23.
@@ E.6 Question Proposal
We generate questions through a brute-force search over various combinations and quantities of conditions. The full set of considered conditions is shown in Table˜1. Specifically, we enumerate all possible attribute combinations containing 0 to 3 conditions and manually select 73 meaningful combinations that can be naturally expressed in language. The selected combinations are listed in Table˜8.
@@ E.7 Question Generation
Given a set of condition types, corresponding values, and a target value, we prompt GPT-4.1 to generate natural language questions using the template shown in Figure˜24.
@@ E.8 Question Quality Control
We employ o4-mini for final quality control. Specifically, o4-mini is used to rank target candidates under two settings: (1) without any supporting context, and (2) with supporting context provided in the form of either summaries or raw conversations, using the prompting format shown in Figure˜25.For each instance, we compute the instance-wise NDCG@10 score in the no-context setting, denoted as , and define the contextual score as , where and are scores under raw and summarized contexts, respectively.